# Thinking Before Running! Efficient Code Generation with Thorough Exploration and Optimal Refinement

### v1

@@ Abstract
Code generation is crucial in software engineering for automating the coding process efficiently. While test-time computation methods show promise, they suffer from high latency due to multiple computation rounds. To overcome this, we introduce ThinkCoder, a framework that combines thorough exploration with optimal refinement. The exploration phase diversifies the solution space by searching for potential solutions, followed by a refinement phase that enhances precision. This approach allows us to select the best solution through careful consideration before taking action, avoiding excessive trial and error. To further minimize test-time computation overhead, we introduce preference-driven optimization with Reinforced Self-Training (ReST), which uses exploration trajectories from ThinkCoder to guide LLM’s evolution. By learning preferences, this approach improves LLM’s exploration efficiency, reducing computational costs while maintaining accuracy. ThinkCoder boosts the performance of multiple base LLMs, excelling on benchmarks like HumanEval and MBPP. Compared to SOTA models, it improves Pass@1 by 1.5% over MapCoder with just 21.7% of the computation cost. Against AgentCoder, ThinkCoder achieves a 0.6% higher Pass@1 after 2 rounds, outperforming AgentCoder’s 5 rounds. Additionally, ReST with success trajectories enhances efficiency, allowing models like LLaMA2-7B to achieve competitive results using only 20% of the computational resources. These results highlight the framework’s effectiveness and scalability. The code will be released upon acceptance.
@@
Xiaoqing Zhang1   Yuhan  Liu1   Flood Sung2  Xiuying Chen311footnotemark: 1     Rui  Yan1thanks:   Corresponding authors. 1Gaoling School of Artificial Intelligence, Renmin University of China  2MoonshotAI 3Mohamed bin Zayed University of Artificial Intelligence {xiaoqingz, yuhan.liu, ruiyan}@ruc.edu.cn {floodsung}@moonshot.cn {xy-chen}@pku.edu.cn
@@ 1 Introduction
Recent research advances indicate that large language models (LLMs) have demonstrated remarkable capabilities in various programming-related domains, such as code generation Zheng et al. (2023); Chaudhary (2023); Dong et al. (2024); Li et al. (2024a), code refinement Chen et al. (2023); Guo et al. (2024); Zheng et al. (2024); Ridnik et al. (2024); Liu et al. (2024); Zhang et al. (2024), and software testing Li et al. (2023); Jalil et al. (2023); Wang et al. (2024); Jones et al. (2024). By utilizing extensive training data, LLMs can understand complex programming tasks, generate syntactically accurate code, and even enhance the quality of their solutions through iterative improvement.
@@
In the context of code generation, researchers have made significant progress in improving prompts to encourage LLMs to generate higher-quality answers. This includes two main directions: optimizing the path for answer exploration and incorporating rich reflection information. In the first direction of path optimization, for example, Self-Planning Jiang et al. (2024) reduces problem complexity in code generation by introducing a planning phase, which refines the process of solution generation. Similarly, CodeCoT Huang et al. (2023a) improves the exploration by employing a Chain-of-Thought prompt, enabling models to reason through each step systematically. In the second direction, the integration of reflection, test cases, and feedback is progressively being used to steer and enhance the generation process. For instance, RethinkMCTS Li et al. (2024b) integrates verbal feedback into its Monte Carlo Tree Search framework to correct errors and improve code generation. Similarly, AgentCoder Huang et al. (2023b) boosts efficiency and accuracy by using interactions between multiple agents to provide valuable input.
Despite their impressive capabilities, a critical challenge remains in fully exploiting the potential of LLMs for high-quality code generation, as their performance is still constrained by issues such as high computation costs and limited effectiveness in integrating iterative feedback. To enhance efficiency and minimize trial-and-error costs, it’s essential to define the improvement direction clearly before implementing refinement. While existing methods like MapCoder, which leverages four LLM-based agents, consumes a considerable amount of computational resources.
@@
To address these challenges, we propose ThinkCoder, a novel framework for efficient code generation that combines thorough exploration and optimal refinement. First, we simplified our multi-agent framework, which now consists only of the Exploration Agent and the Execution Agent. The Exploration Agent is responsible for deep thinking and expanding exploration with various codes and tests. The Execution Agent independently verifies code during the optimal refinement process, without relying on LLMs. By utilizing a dynamically created testing pool, ThinkCoder ensures accurate verification of the generated code and offers reliable guidance for further optimization. Second, to tackle the additional computational load resulting from extensive exploration, we incorporate a preference-driven optimization phase using Reinforced Self-Training (ReST) in ThinkCoder. ReST leverages successful exploration trajectories to train LLM, enabling it to generate optimal solutions more efficiently during exploration. This approach significantly reduces the computational cost of test-time computations, enabling models like LLaMA2-7B to achieve notable improvements in both performance and efficiency.
@@
$\bullet$ We propose ThinkCoder, an efficiency framework for code generation that incorporates only two agents: an exploration agent and an execution agent that enables thorough exploration and optimal refinement, ensuring high-quality code generation.
$\bullet$ We introduce preference-driven optimization with ReST, using exploration trajectories from ThinkCoder to enhance the generation and verification capabilities of LLMs, further reducing computational overhead during testing.
@@ 2 Related Work
Multi-Agent Code Generation. Deriving multiple agents by mimicking the different stages of human programming can effectively improve both the quality and efficiency of code generation. For instance, AgentCoder Huang et al. (2023b) introduces programmer agents, test design agents, and test execution agents, integrating code generation, test case generation, and test feedback into the code generation process. Similarly, MapCoder Islam et al. (2024) introduces retrieval agents, planning agents, coding agents, and debugging agents, iterating on code generation quality through more frequent agent communication and reflection. However, MapCoder incurs higher computational overhead and relies solely on sample I/O for code execution, limiting test case coverage and robustness. In contrast, our framework, ThinkCoder, addresses these limitations by self-verification efficiently without excessive computational costs.
Figure 1: The end-to-end process of ThinkCoder involves $k$ thorough exploration steps followed by $n$ optimal refinement cycles. The Exploration Agent generates $k$ codes and tests simultaneously, storing results in a Testing Pool. The optimal refinement phase includes self-verification that selects the optimal solution with the Execution Agent and aggregate reflection for the instruction update. The optimal refinement will be repeated recursively for $n$ cycles, ultimately leading to the final solution.
@@ 3.1 Overall Architecture
Figure 1 illustrates the structure of ThinkCoder. Given an instruction $I=[p,\text{prompt}]$ for problem $p$, where prompt is the task description for code generation, the Exploration Agent produces $k$ candidate solutions. Each solution is paired with $m$ test cases, and the temperature $t$ of the LLM’s generation is adjusted to promote diversity. These test cases are then aggregated into a Testing Pool $TP$. We define the exploration results as $E=[G,TP]$, where $G=\{g_{i}\mid g_{i}=\text{LLM}(p,t)\text{ for }i\in\{1,2,\dots,k\}\}$, and $TP\in\mathbb{T}^{k\times m}$, where $\mathbb{T}$ represents each test. The Execution Agent sequentially executes each candidate solution in $G$ using the test cases from $TP$ to compute pass rates. The code $g_{s}$ with the highest pass rate is selected for further refinement. Here we define the pass rate of $g_{i}$ as $r_{g_{i}}=\frac{\text{number of tests passed by }g_{i}}{|TP|}$ and $g_{s}=\arg\max_{i}r_{g_{i}}$. Feedback $f$ from failed test case $f_{t}$, along with the current code $g_{s}$, generated test cases $tests$, and problem $p$, is used to update the instructions $I=[p,g_{s},f_{t},f,tests,prompt]$ for the next exploration. This iterative process allows the Exploration Agent to generate improved solutions and test cases over $n$ iterations. To avoid unnecessary computation on well-solved problems, if $r_{g_{i}}$ is greater than $\theta$, the refinement process is terminated, and the code $g_{s}$ is returned as the final solution. Please refer to the ThinkCoder algorithm 1 for more detailed information on these operations. Below, we detail the implementation of Thorough Exploration, Optimal Refinement, Exploration&Execution Agent, and Testing Pool.
Algorithm 1 ThinkCoder
1:Initialize the testing pool TP=[], set the globally max pass rate $g_{p}=0$ and best solution $g_{s}=None$
2:Set the hyperparameters $t=0.5,k=5,n=5,m=3,\theta=0.8$
3:Set the instruction $prompt$ for code generation task and instruction for task $p$ as $I=[p,prompt]$
4:for each $iter\in n$ do
5:     Initialize the locally max pass rate $l_{p}=0$ and best solution ${l_{s}}={None}$ for iteration $iter$
6:     Randomly select $m$ successful test cases from TP as $tests$
7:     Select a new type of error feedback for test $f_{t}$ from TP as $f$
8:     if $f_{t}$ and $f$ is not None then
9:         Update the instruction $I=[p,g_{s},f_{t},f,tests,prompt]$
10:     end if
11:     Get code solutions $G^{\prime}$ and test cases $TP^{\prime}$ from exploration agent for the instruction $I$ within $k$ iterations
12:     Deduplicate testing pool TP’ by AST
13:     for each $i\in|G^{\prime}|$ do
14:         Get the pass rate $r_{g_{i}}$ for solution $g_{i}$ from execution agent with the input $TP^{\prime}$
15:         if $r_{g_{i}}>l_{p}$ then
16:              Set $l_{p}=r_{g_{i}},l_{s}={g_{i}}$
17:         end if
18:     end for
19:     if $l_{p}>g_{p}$ then
20:         Set $g_{p}=l_{p},g_{s}=l_{s}$
21:         Add TP’ to TP
22:     end if
23:     if $g_{p}>\theta$ then
24:         return $g_{s}$ as the best solution
25:         Terminate optimal refinement iteration
26:     end if
27:end for
@@ 3.2 Thorough Exploration
Before running the generated code and performing corrections, ThinkCoder conducted extensive thinking during the Thorough Exploration phase. By generating $k$ answers at once, it expanded the search area. The search for multiple answers was achieved by increasing the LLM’s temperature. With the increasing temperature, the probability distribution of the LLM becomes flatter, reducing the differences between word probabilities and increasing the entropy of the distribution, which enhances the diversity of the generated text. The first iteration of exploration aims to generate diverse code and comprehensive test cases, including regular tests, boundary tests, and performance tests. While the subsequent iterations, focus on the exploration of code correction for specific errors. The exploration is carried out by the Exploration Agent, which is the only LLM used in ThinkCoder. The prompt design refers to Appendix A.3.
@@ 3.3 Optimal Refinement
To solve the problem efficiently, it is essential to first identify the optimal solution through careful consideration before making any code corrections. Therefore, Optimal Refinement involves selecting the best solution via self-verification and gathering the necessary information for deeper exploration and improvement in the next exploration. During self-verification, code quality is evaluated by the pass rate, determined by the Execution Agent running candidate code against all test cases. The code with the highest pass rate is selected for further refinement. From this, we randomly select an unresolved error to target for the current round of improvements. This ensures all errors are eventually addressed. The refinement instruction is created using the following information: the problem, the best solution identified so far, feedback, and test cases. After each refinement, the solution’s pass rate is evaluated. If it surpasses any previous solution, it is marked as the best. Through multiple iterations, the best solution is continuously corrected and optimized, leading to a gradual increase in the pass rate within the testing pool. For more details, please refer to the Appendix A.3.
3.4 Exploration & Execution Agent
ThinkAgent consists of two agents: the exploration agent and the execution agent. The exploration agent uses LLMs to search for diversity code and test cases, while the execution agent runs the provided code snippets and test cases in a local environment to generate feedback. To improve the efficiency of the execution agent, we leverage multiprocessing to validate the code in parallel.
@@ 3.5 Testing Pool
The Testing Pool contains the test cases generated by the exploration agent which contains significantly more test cases than the ground truth tests. These test cases are categorized into three types: regular testing, boundary testing, and performance testing. A higher pass rate in the testing pool reflects stronger code robustness, making it essential to ensure the diversity and accuracy of the test cases. Several key operations contribute to achieving this. Abstract Syntax Tree (AST)  Backus et al. (1960) is applied to identify and remove duplicate test cases in the testing pool, eliminating unnecessary time costs associated with redundant tests. Additionally, after each iteration, the testing pool only selects test cases that can improve the code’s pass rate, thus enhancing the diversity and accuracy of the pool. Finally, the testing pool will select test cases with distinct errors and use feedback to update the code’s improvement goals, driving continuous optimization.
@@ 4.1 Data Collection
Temperature-based Generation. During temperature-based generation, we obtain $D_{M_{0}}$ to enhance the diversity of responses. For a given problem $x$, the language model $\mathcal{M}_{0}$ is used to explore $k$ times, generating $\{y^{j}\}_{j=1}^{k}$, where $y^{j}\sim\mathcal{M}_{0}(y|x)$. During exploration, a relatively large temperature value $t$ is set for $\mathcal{M}_{0}$. The generated outputs are then evaluated using the ExecuteAgent, which computes $E(x,y^{j})=(pass\_rate,feedback)$. For any sample where $pass\_rate==1$, the corresponding instance $(x,y^{j})$ is added to the training dataset $D_{M_{0}}$.
Reflection-based Generation. The primary purpose of generating $R_{M_{0}}$ is to enhance the ability of error correction. For a given sample $x$ and the feedback $E(x,y^{j})=(pass\_rate,feedback)$ from the ExecuteAgent, for samples with $pass\_rate=0$, we add the feedback to $x$ and use the language model $\mathcal{M}_{0}$ to perform temperature-based exploration on the input $x^{\prime}=(x,feedback)$ and get the output $y^{\prime j}$. Then we also use the execution agent to obtain the right correction answer with $E(x^{\prime},y^{\prime j})=(pass\_rate^{\prime},feedback^{\prime})$. Finally, we add the instance $(x^{\prime},y^{\prime j})$ with the $pass\_rate^{\prime}==1$ to the reflection-based trajectories $R_{M_{0}}$.

### v2

@@ Abstract
Code generation is crucial in software engineering for automating the coding process efficiently. While test-time computation methods show promise, they suffer from high latency due to multiple computation rounds. To overcome this, we introduce ThinkCoder, a framework that combines thorough exploration with optimal refinement. The exploration phase diversifies the solution space by searching for potential solutions, followed by a refinement phase that enhances precision. This approach allows us to select the best solution through careful consideration before taking action, avoiding excessive trial and error. To further minimize test-time computation overhead, we introduce preference-driven optimization with Reinforced Self-Training (ReST), which uses exploration trajectories from ThinkCoder to guide LLM’s evolution. This approach enhances LLM’s exploration efficiency via preference learning, cutting costs while maintaining accuracy. ThinkCoder boosts the performance with a single LLM, excelling on benchmarks like HumanEval and MBPP. Compared to SOTA models, it improves Pass@1 by 3.0% over MapCoder with just 6.4% of the computation cost. Against AgentCoder, ThinkCoder achieves a 0.5% higher Pass@1 after 2 rounds, outperforming AgentCoder’s 5 rounds. Additionally, ReST with success trajectories enhances efficiency, allowing models like LLaMA2-7B to achieve competitive results using only 20% of the computational resources. These results highlight the framework’s effectiveness and scalability.  [^1] The code is available at https://github.com/xiaoqzhwhu/ThinkCoder
@@
Xiaoqing Zhang1,2thanks: This work was done during the internship at Moonshot AI.   Yuhan  Liu122footnotemark: 2   Flood Sung2  Xiuying Chen3     Shuo  Shang4   Rui  Yan1,5,6thanks:   Corresponding authors. 1Gaoling School of Artificial Intelligence, Renmin University of China  2Moonshot AI 3Mohamed bin Zayed University of Artificial Intelligence 4University of Electronic Science and Technology of China  5School of Artifcial Intelligence, Wuhan University 6Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MoE {xiaoqingz, yuhan.liu, ruiyan}@ruc.edu.cnfloodsung@moonshot.cn xy-chen@pku.edu.cnjedi.shang@gmail.com
@@ 1 Introduction
Recent research advances indicate that large language models (LLMs) have demonstrated remarkable capabilities in various programming-related domains, such as code generation Zheng et al. (2023); Chaudhary (2023); Dong et al. (2024); Li et al. (2024a), code refinement Chen et al. (2023); Guo et al. (2024); Zheng et al. (2024); Ridnik et al. (2024); Liu et al. (2024a); Zhang et al. (2024), and software testing Li et al. (2023); Jalil et al. (2023); Wang et al. (2024); Jones et al. (2024). By utilizing extensive training data, LLMs can understand complex programming tasks, generate syntactically accurate code, and even enhance the quality of their solutions through iterative improvement.
@@
In the context of code generation, researchers have made significant progress in improving prompts to encourage LLMs to generate higher-quality answers. This includes two main directions: optimizing the path for answer exploration and incorporating rich reflection information. In the first direction of path optimization, for example, Self-Planning Jiang et al. (2024) reduces problem complexity in code generation by introducing a planning phase, which refines the solution generation process. Similarly, CodeCoT Huang et al. (2023a) improves the exploration by employing a Chain-of-Thought(CoT) prompt, enabling models to reason through each step systematically. In the second direction, the integration of reflection, test cases, and feedback is progressively being used to steer and enhance the generation process. For instance, RethinkMCTS Li et al. (2024b) integrates verbal feedback into its Monte Carlo Tree Search framework to correct errors and improve code generation. Similarly, AgentCoder Huang et al. (2023b) boosts efficiency and accuracy by using interactions between multiple agents to provide valuable input.
Despite the significant potential of using CoT for planning code improvements, the thought process incurs a substantial token overhead. To enhance efficiency and minimize trial-and-error costs, it is crucial to clarify the optimization direction before implementing refinement. Existing methods, such as MapCoder, which leverages four LLM-based agents, consume a considerable amount of computational resources.
Based on these insights, we propose ThinkCoder, a novel framework for efficient code generation that combines thorough exploration and optimal refinement. First, we simplified the multi-agent framework into a single Exploration Agent and a CodeVerifier. The Exploration Agent is responsible for in-depth analysis and generating various codes, while exploring diverse test cases to build a robust testing pool. The CodeVerifier, using the testing pool, independently determines the optimization direction without relying on LLMs. By continuously improving the testing pool throughout the refinement process, ThinkCoder ensures accurate verification of generated code in each round, providing efficient guidance for subsequent optimizations.
@@
Second, to tackle the additional computational load resulting from extensive exploration, we incorporate a preference-driven optimization phase using Reinforced Self-Training (ReST) in ThinkCoder. ReST leverages successful exploration trajectories to train LLM, enabling it to generate optimal solutions more efficiently during exploration. This approach significantly reduces the computational cost of test-time computations, enabling models like LLaMA2-7B to achieve notable improvements in both performance and efficiency.
@@
$\bullet$ We introduce ThinkCoder, an efficiency framework for code generation that integrates an LLM-based exploration agent with a non-LLM-based CodeVerifier. We utilize a self-evolving testing pool, significantly enhancing the speed of code verification and improvement.
$\bullet$ We introduce preference-driven optimization with ReST, using exploration trajectories from ThinkCoder to enhance LLMs’ generation and verification capabilities, further reducing computational overhead during testing.
@@ 2 Related Work
Multi-Agent Code Generation. Deriving multiple agents by mimicking the different stages of human programming can effectively improve both the quality and efficiency of code generation, as agents based on large language models possess strong reasoning capabilities Liu et al. (2024b, c); Zhang et al. (2025). For instance, AgentCoder Huang et al. (2023b) introduces programmer agents, test design agents, and test execution agents during the code generation process. Similarly, MapCoder Islam et al. (2024) introduces retrieval agents, planning agents, coding agents, and debugging agents, iterating on code generation quality through more frequent agent communication and reflection. However, MapCoder incurs higher computational overhead and relies solely on sample I/O for code execution, limiting test case coverage and robustness. In contrast, our framework, ThinkCoder, addresses these limitations by self-verification efficiently without excessive computational costs.
Figure 1: The end-to-end process of ThinkCoder involves $k$ thorough exploration steps followed by $n$ optimal refinement cycles. The Exploration Agent generates $k$ codes and tests simultaneously during self-exploration, storing results in a Testing Pool. The optimal refinement phase includes self-verification that selects the optimal solution with the CodeVerifier and aggregate reflection for the instruction update. The optimal refinement will be repeated recursively for $n$ cycles, ultimately leading to the final solution.
@@ 3.1 Overall Architecture
Figure 1 illustrates the structure of ThinkCoder. Given an instruction $I=[p,\text{prompt}]$ for problem $p$, where prompt is the task description for code generation, the Exploration Agent produces $k$ candidate solutions. Each solution is paiblue with $m$ test cases, and the temperature $t$ of the LLM’s generation is adjusted to promote diversity. These test cases are then aggregated into a Testing Pool $TP$. We define the exploration results as $E=[G,TP]$, where $G=\{g_{i}\mid g_{i}=\text{LLM}(p,t)\text{ for }i\in\{1,2,\dots,k\}\}$, and $TP\in\mathbb{T}^{k\times m}$, where $\mathbb{T}$ represents each test. The CodeVerifier sequentially executes each candidate solution in $G$ using the test cases from $TP$ to compute pass rates. The code $g_{s}$ with the highest pass rate is selected for further refinement. Here we define the pass rate of $g_{i}$ as $r_{g_{i}}=\frac{\text{number of tests passed by }g_{i}}{|TP|}$ and $g_{s}=\arg\max_{i}r_{g_{i}}$. Feedback $f$ from failed test case $f_{t}$, along with the current code $g_{s}$, generated test cases $tests$, and problem $p$, is used to update the instructions $I=[p,g_{s},f_{t},f,\text{tests},\text{prompt}]$ for the next exploration. This iterative process allows the Exploration Agent to generate improved solutions and test cases over $n$ iterations. To avoid unnecessary computation on well-solved problems, if $r_{g_{i}}$ is greater than $\theta$, the refinement process is terminated, and the code $g_{s}$ is returned as the final solution. For more details of ThinkCoder please refer to the algorithm 1 in Appendix A.1. Below, we detail the implementation of Thorough Exploration, Optimal Refinement, Exploration Agent&CodeVerifier, and Testing Pool.
@@ 3.2 Thorough Exploration
During the thorough exploration phase, ThinkCoder conducts extensive thinking before running code and performing corrections. The self-exploration process generates $k$ answers simultaneously to expand the search area by increasing the LLM’s temperature, which flattens the probability distribution and enhances diversity. The first iteration of self-exploration aims to generate diverse code and comprehensive test cases, including regular tests, boundary tests, and performance tests. While the subsequent iterations, focus on the exploration of code correction for specific errors. The only LLM-based agent conducts self-exploration with prompt design detailed in Appendix A.4.
@@ 3.3 Optimal Refinement
To solve the problem efficiently, it is essential to first identify the optimal solution through careful consideration before making any code corrections. Therefore, optimal refinement involves selecting the best solution via self-verification and gathering the necessary information for deeper exploration and improvement in the next exploration. During self-verification, code quality is assessed by the pass rate, calculated by the CodeVerifier running candidate code against all test cases. The code with the highest pass rate is chosen for refinement. We then randomly select an unresolved error to target in the current round of improvements, ensuring all errors are eventually addressed. The refinement instruction is generated based on the problem, the current best solution, feedback, and test cases. After each refinement, the solution’s pass rate is evaluated. If it exceeds previous solutions, it becomes the new best. Through multiple iterations, the best solution is continuously corrected and optimized, resulting in a gradual increase in the pass rate within the testing pool. For more details, please refer to the Appendix A.4.
3.4 Exploration Agent& CodeVerifier
ThinkAgent consists of a single exploration agent and a CodeVerifier. The exploration agent uses LLMs to search for diversity code and test cases, while the CodeVerifier runs the provided code snippets and test cases in a local environment to generate feedback. To improve the efficiency of the CodeVerifier, we leverage multiprocessing to validate the code in parallel.
@@ 3.5 Testing Pool
The Testing Pool includes numerous test cases generated by the exploration agent, categorized into regular, boundary, and performance tests. Our goal is to create a robust pool to assess complex situations beyond simple cases. A higher pass rate indicates stronger code robustness, so we ensure test case diversity and accuracy. Several key operations contribute to achieving this. Abstract Syntax Tree (AST)  Backus et al. (1960) is applied to identify and remove duplicate test cases in the testing pool, eliminating unnecessary time costs associated with redundant tests. Additionally, after each iteration, the testing pool only selects test cases that can improve the code’s pass rate, thus enhancing the diversity and accuracy of the pool. Finally, the testing pool will select test cases with distinct errors and use feedback to update the code’s improvement goals, driving continuous optimization.
4 Preference-driven Optimization
@@ 4.1 Data Collection
[…]
GPT-3.5-Turbo MapCoder 78.3 54.4 80.5 70.1 70.8
AgentCoder 89.9 89.1 79.9 77.4 84.1
ThinkCoder(Ours) 90.9 78.8 89.0 81.0 84.9
GPT-4-Turbo MapCoder 83.1 57.7 93.9 82.9 79.4
AgentCoder 91.4 91.4 89.6 76.2 87.2
ThinkCoder(Ours) 90.5 80.5 93.9 86.0 87.7
Table 2: Pass@1 for MapCoder, AgentCoder, and ThinkCoder on MBPP, MBPP-ET, HumanEval, and HumanEval-ET benchmarks with GPT-3.5-Turbo and GPT-4-Turbo as the backbone.
[…]
GPT-4o LDB 82.4 65.4 98.2 81.7 29.3 71.4
LPW 84.8 65.8 98.2 84.8 34.7 73.7
ThinkCoder(Ours) 90.0 79.2 97.0 89.6 40.1 79.2
Table 3: Pass@1 for LDB, LPW, and ThinkCoder on MBPP, MBPP-ET, HumanEval, HumanEval-ET, and CodeContests benchmarks with GPT-4o as the backbone.